OpenStudy (anonymous):

HELP ME PLEASE THANKS Match each percent proportion with the percent problem it represents TERM definition 80/35=x100 A-what is 80% of 35? x/35=80/100 B-35 is 80% of what? 80/100=35/x C-80 is what percent of 35? 35/100=80/x D-35% of what number is 80?

OpenStudy (anonymous):

anyone please

OpenStudy (anonymous):

thanks @eSpeX for spending your time looking at the question

OpenStudy (espex):

It is a bit difficult with the formatting offered on these sites, let me see if I can clean it up for clarification.

OpenStudy (anonymous):

okay

OpenStudy (espex):

If I am reading it correctly your terms are: 80/35 = x% x/35 = 80% 80/100 = 35/x 35/100 = 80/x Is that right?

OpenStudy (anonymous):

80/35 = x/100 x/35 = 80/100 80/100 = 35/x 35/100 = 80/x

OpenStudy (anonymous):

i wrote it wrong the first time, i'm sorry

OpenStudy (espex):

You know that a fraction is just a representation of a "part" of something, right? And a percentage is also a "part" of something. So taking the first line: 80/35=x/100 and read it as 80 parts of 35 equals how much of the whole (100). Had to correct for the terms. Make sense?

OpenStudy (anonymous):

yea now it makes sence

OpenStudy (espex):

That's great. If you need more help don't hesitate to ask. :)